Alex\ is\ building\ a\ rectangular\ garden\ that\ has\ an\ area\ of\ 912\ square\ feet.\ His\ wife\ decided\ that\ she\ wants\ the\ rectangular\ garden\ to\ have\ an\ area\ of\ 972\ square\ feet.\ Alex\ can\ make\ the\ change\ if\ he\ reduces\ the\ width\ of\ the\ garden\ that\ he\ was\ planning\ by\ 2\ feet\ and\ increase\ the\ length\ of\ the\ garden\ that\ he\ was\ planning\ by\ 3\ feet.\ What\ is\ the\ dimensions\ of\ the\ garden\ that\ Alex\ wife\ wants?
Found 2 solutions by josgarithmetic, Alan3354:
Answer by josgarithmetic(39838)

\\\\\\\\\\........

Answer by Alan3354(69443)

Alex\ is\ building\ a\ rectangular\ garden\ that\ has\ an\ area\ of\ 912\ square\ feet.\ His\ wife\ decided\ that\ she\ wants\ the\ rectangular\ garden\ to\ have\ an\ area\ of\ 972\ square\ feet.\ Alex\ can\ make\ the\ change\ if\ he\ reduces\ the\ width\ of\ the\ garden\ that\ he\ was\ planning\ by\ 2\ feet\ and\ increase\ the\ length\ of\ the\ garden\ that\ he\ was\ planning\ by\ 3\ feet.\ What\ is\ the\ dimensions\ of\ the\ garden\ that\ Alex\ wife\ wants?
------------
L*W = 912
(L+3)*(W-2) = 972
----
L = 912/W
Sub for L
---
(912/W + 3)*(W-2) = 972
912 - (1824/W) + 3W - 6 = 972
912W - 1824 + 3W^2 - 6W = 972W
3W^2 - 66W - 1824 = 0
Can you do the rest?
=================
= = = = = = = = =
You need a space bar.
If you join T Rump's Space Force, you can go to space bars.